OpenStudy (anonymous):

Find the equation of the line through (4,3) perpendicular to the line 3x+y=7

OpenStudy (solomonzelman):

Put it in y=mx+b form, y=-3x+7 perpendicular, that means that the slope is going to be a multiplicative inverse times '-1'. For slope b, it's -1/b So here it's y=(1/3)x+b we don't know the y=intercept yet, so plug in your point, (3)=(1/3)(4)+b 3=4/3+b 3-4/3=4/3-4/3+b 5/3=b So, y=(1/3)x+5/3

OpenStudy (anonymous):

how did you get the slope as being 1/3?

OpenStudy (solomonzelman):

because a perpendicular slope for m, is -1/m (multiplicative inverse and then times '-1') so the multiplicative inverse of -3 is -1/3 and times -1 is 1/3

OpenStudy (anonymous):

thank you!

OpenStudy (solomonzelman):

Anytime!